OpenStudy (anonymous):

if f(0)=k f(x)=(x^2-x)/(2x) and f is discontinuous at x=0, then k = A)-1 B)-1/s C)0 D)1/s E)1 can someone explain how to solve this?

OpenStudy (freckles):

So wait a minute are you saying f(x)= k if x=0 and (x^2-x)/(2x) if x doesn't equal 0 and you have the question if f is discontinuous at x=0, then k=?

OpenStudy (freckles):

what is s?

OpenStudy (freckles):

is s, two?

OpenStudy (freckles):

And are you sure we don't want to find k so that f is continuous at x=0?

OpenStudy (freckles):

Because that question would make more sense?

OpenStudy (anonymous):

uhm let me send a picture of what it looks like

OpenStudy (freckles):

trick is to try to reduce that one fraction first

OpenStudy (anonymous):

so wouldn't it be x(x-1)/2x?

OpenStudy (freckles):

then cancel a factor x on top and bottom

OpenStudy (freckles):

this is what we want \[\lim_{x \rightarrow 0}\frac{x-1}{2} =k\]

OpenStudy (freckles):

recall that we want \[\lim_{x \rightarrow 0}f(x)=f(0) \] if we have that then it is continuous at x=0

OpenStudy (unklerhaukus):

s=2

OpenStudy (freckles):

where is the dolphin? Are you still there?

OpenStudy (anonymous):

yes I'm here

OpenStudy (anonymous):

so then how would we find k?

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{x-1}{2} =k\] Does this equation make any sense to you?

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}f(x)=f(0) \\ \lim_{x \rightarrow 0}\frac{x-1}{2}=f(0) \\ \lim_{x \rightarrow 0}\frac{x-1}{2}=k\]

OpenStudy (freckles):

find that limit and you are done because k is equal to what limit you get there

OpenStudy (anonymous):

oh so then i would just plug in zero and get -1/2 right?

OpenStudy (freckles):

yea

OpenStudy (freckles):

the function (x-1)/2 is continuous at x=0 so you can just plug in 0 for x

OpenStudy (freckles):

that is how uncle and i already knew what you meant by s

OpenStudy (freckles):

lol

OpenStudy (freckles):

s had to be two in order for there to be an answer

OpenStudy (anonymous):

oh lol okay. I see. well thank you very much. You helped lots !

OpenStudy (freckles):

np